Check the convergence of the sequence defined by 𝑒𝑛+1 = 1 2 (𝑒𝑛 + π‘Ž 𝑒𝑛 ) , π‘Ž > 0. Note that this is the sequence associated with finding the square root of a number π‘Ž > 0 by the Newton’s method.


1
Expert's answer
2021-11-04T17:15:55-0400

let the sequenceΒ "u_n"Β converges toΒ "l"Β .

Then,

"u_{n+1}=\frac{1}{2}(u_n+\frac{a}{u_n})\\
l=\frac{1}{2}(l+\frac{a}{l})\\
\text{Solving this equation,we get}\\
l=\frac{1}{2}(\frac{a+l^2}{l})\\
2l^2=a+l^2\\
l^2=a\\
l=\pm\sqrt{a}\\
\text{Since, it is given a>0. Therefore, }\\l=\sqrt{a}"


Checking hypothesis that the sequence converges, we get the value of sequence limit "\sqrt{a}"

So, we can conclude that this sequence converges.